# Implicit In-context Learning

### v1

@@ 4.1 Benchmarking I2CL
Table 1: Comparison between I2CL and other methods on Llama2-7b. Mean accuracy and standard deviation across five random seeds are reported for each task. The best results are highlighted in bold, and the second-best results are shown in blue. In addition to a practical gauge of the inference speed and memory usage (see Fig 1), we include an examination of cached parameters. Here, $M$, $D$, and $L$ denote the number of demonstration tokens, model dimension, and architecture layers, respectively. $P$ indicates the number of extra learnable tokens in the Soft-prompt method, and $1/K$ represents the compression rate in the Label-anchor method.
@@
Task Zero-shot Few-shot (ICL) Noise vector Soft-prompt Label-anchor Task-vector I2CL (ours) SST-2 (%) $\uparrow$ $83.00$ $94.44$$\pm$1.44 $49.88$$\pm$0.24 $56.24$$\pm$6.99 $83.32$$\pm$5.95 $81.44$$\pm$4.73 $\mathbf{87.68}$$\pm$2.47 SST-5 (%) $\uparrow$ $27.00$ $41.72$$\pm$3.68 $20.56$$\pm$0.64 $24.24$$\pm$2.96 $27.68$$\pm$4.21 $25.96$$\pm$0.59 $\mathbf{39.12}$$\pm$2.69 TREC (%) $\uparrow$ $50.00$ $77.32$$\pm$4.41 $20.12$$\pm$10.92 $55.20$$\pm$4.14 $77.48$$\pm$3.49 $65.68$$\pm$1.93 $\mathbf{78.56}$$\pm$5.32 AGNews (%) $\uparrow$ $70.20$ $85.68$$\pm$2.00 $27.32$$\pm$2.82 $78.00$$\pm$7.60 $83.72$$\pm$1.04 $79.68$$\pm$4.07 $\mathbf{85.48}$$\pm$1.16 Subj (%) $\uparrow$ $51.40$ $52.56$$\pm$3.09 $49.64$$\pm$0.48 $57.40$$\pm$4.93 $53.00$$\pm$2.95 $58.56$$\pm$4.91 $\mathbf{73.84}$$\pm$3.84 HateSpeech18 (%) $\uparrow$ $54.20$ $70.24$$\pm$5.80 $59.84$$\pm$8.04 $59.56$$\pm$6.96 $64.52$$\pm$8.09 $67.68$$\pm$3.70 $\mathbf{69.88}$$\pm$5.67 DBPedia (%) $\uparrow$ $72.00$ $96.64$$\pm$0.48 $7.28$$\pm$0.37 $74.40$$\pm$6.43 $81.40$$\pm$3.67 $89.48$$\pm$2.58 $\mathbf{90.16}$$\pm$1.86 EmoC (%) $\uparrow$ $41.80$ $75.48$$\pm$1.63 $26.76$$\pm$3.04 $35.08$$\pm$5.29 $59.12$$\pm$10.60 $44.64$$\pm$3.53 $\mathbf{63.72}$$\pm$1.37 MR (%) $\uparrow$ $73.60$ $93.24$$\pm$0.50 $50.12$$\pm$0.24 $54.32$$\pm$1.76 $84.40$$\pm$5.89 $82.32$$\pm$5.37 $\mathbf{87.68}$$\pm$2.26 Macro avg. acc. (%) $\uparrow$ $58.13$ $76.37$ $34.61$ $54.94$ $68.29$ $66.16$ $\mathbf{75.12}$#\## cached parameters $\downarrow$ 00 $2MDL$ $2DL$ $(2M+P)DL$ $2(M/K)DL$ $D$ $2DL$
I2CL Achieves Few-shot Performance at Zero-shot Cost I2CL is designed to enhance zero-shot performance, aspiring to match or exceed those of few-shot outcomes. As shown in Table 1, I2CL significantly outperforms the zero-shot counterpart by 17% in absolute accuracy and is marginally behind (by around 1%) the few-shot learning. An interesting outcome arises from the Subj task, where the instructions (i.e., demonstration examples) do not function effectively with ICL, yet are well adhered to under I2CL. Critically, I2CL caches only a fixed number of vectors, independent of the context token length, and retains the speed of zero-shot learning. These results demonstrate I2CL’s potential as a more economical alternative to traditional ICL.
Comparison with Other Methods To further validate the effectiveness of I2CL, we compare it with other comparable techniques. (1) Noise vector: We substitute the context vector with random noise in order to assess its necessity. (2) Soft-prompt: Another approach of using demonstration examples is to treat them as training data and apply the soft-prompt technique. To this end, we compare I2CL with representative Prompt-tuning [Lester et al., 2021] method to demonstrate the superior data efficiency of I2CL. (3) Label-anchor: We also compare I2CL with the context compression technique presented in Wang et al. [2023] under the same setup. Following the original methodology, we include formatting tokens as anchors. (4) Task-vector: We compare I2CL to the task vector [Hendel et al., 2023] method, which also improves zero-shot performance. Adhering to their setup, we use a hold-out validation set to identify the optimal replacement layer for each task. We refer to Appendix B for reproduction details. As displayed in Table 1, replacing the context vector with random noise significantly degrades performance, yielding subpar results compared to zero-shot. The soft-prompt technique is beneficial for only a few tasks under data scarcity, and it can also impair the performance. The label-anchor method exhibits a decent upgrade over the zero-shot baseline, achieving results most close to I2CL. However, its inference cost remains dependent on the length of demonstration tokens, with reductions proportional to their compression rate. Alike I2CL, the task-vector method also enjoys zero-shot inference expense; Nevertheless, its performance is sensitive to the downstream task and is slightly worse on average than the label-anchor method. It also requires an additional hyperparameter selection step, i.e., choosing the target replacement layer using a hold-out validation set. I2CL achieves the best performance on all tasks and surpasses its rivals by a considerable margin.
Figure 3: Left: Evaluation of I2CL and few-shot learning under deficient demonstrations. The symbol * denotes the results under deficient demonstration examples. “Unseen demo” refers to the evaluation of calibrated coefficients on unseen demonstrations. Middle: Analysis of the influencing factors of context vectors. “Random-labe” indicates random input-label mappings. “Random-order” refers to the random permutation of words. “W/o format” signifies excluding the template tokens in the creation of context vectors. Right: t-SNE plot of context vectors. Each circle denotes a context vector generated using a group of randomly sampled demonstration examples.
@@ 4.2 On the Formation of the Context Vector
Deficient Demonstrations $\bm{\neq}$ Deficient I2CL It is well-known that ICL is sensitive to the choice of demonstration examples [Dong et al., 2023], even the order [Zhao et al., 2021, Lu et al., 2022]. In this context, we investigate whether deficient demonstration examples necessarily lead to degraded I2CL performance. We sample $20$ different groups of demonstration examples and keep the group with the poorest ICL performance on a hold-out validation set. Using this poorest group, we then perform both ICL and I2CL on an evaluation set. Referencing Figure 3 (left), the deficient demonstration group leads to a severely downgraded ICL performance ($-7\%$), while I2CL performance is barely affected ($-0.5\%$). This result suggests that less attention can be paid to the selection of demonstration examples when using I2CL. We hypothesize that this resiliency is due to the emergence of more abstract and generic concepts within the context vector, making them robust against token space variations. To corroborate this hypothesis, we visualize the context vectors using the t-SNE technique [van der Maaten and Hinton, 2008] in Figure 3 (right). Context vectors under different demonstration groups are close to each other and different tasks hold distinct context vectors.
@@
Influencing Factors of the Context Vector Generation We have shown that context vectors are robust against the selection of demonstration examples. Here, we explore various factors that may impact the context vector’s functionality. Inspired by counter-intuitive findings reported in [Zhang et al., 2022, Min et al., 2022b], we experiment with two variations of demonstration examples. Random-label: pairing each input sentence with a random label from the task, rather than the ground-truth label. Random-token: randomly permuting all words within a demonstration example. Another ineligible factor involves the inclusion of formatting tokens. W/o-format: removing formatting words from a demonstration example, retaining only the input sentence and its corresponding label, e.g., “Review: This is a great movie. Label: positive. As revealed in Figure 3 (middle), input-label mapping relations have minimal impact on context vector formation, mirroring the observations made in ICL. However, these mapping relations are crucial for calibration purposes; using random input-label mapping during calibration undermines the functionality of I2CL. Unlike the phenomenon observed in [Zhang et al., 2022], word sequence holds essential statistics under a casual architecture. Randomly permuting input sequences yields degenerated context vectors and a clearly downgraded I2CL performance. Lastly, formatting tokens contribute substantially—removing them leads to a noticeable performance drop.
@@ 4.3 Analysis of Calibrated Linear Coefficients
With a grasp on the formation of context vectors, we now delve into the properties of calibrated linear coefficients to enhance our understanding of the internal mechanisms of I2CL.
@@
Linear Coefficients are Generalizable Given the demonstration-dependent context vectorization and calibration process, it is natural to consider I2CL as a test-time optimization framework. However, we show that the coefficients are generalizable and require calibration once per task. In this context, we evaluate a set of calibrated coefficients on five new context vectors generated using five unseen groups of demonstrations. As exhibited in Figure 3 (left), the calibrated linear coefficients generalize well to unseen demonstrations. We attribute this generalization capability to two factors: (1) the inherent robustness of the context vector against token space variations, as demonstrated in Section 4.2, and (2) the sufficiency of calibrated linear coefficients, independent of context vectors and query activation, to uniquely represent a task. We substantiate the second premise in the following subsection.
@@
Calibrated Coefficients Embed Task Semantics Here, we investigate whether calibrated coefficients alone are adequate to serve as "task id". In pursuit of this goal, we concatenate calibrated coefficients across layers to form a 1D vector, which we then visualize using t-SNE. As illustrated in Figure 4 (left), the calibrated linear coefficients are tightly clustered for instances associated with the same task and separate otherwise, providing complementary evidence for their good generalization capability. One exception comes from the proximity between MR and SST-2. This phenomenon is not unexpected, as both tasks originate from the same underlying distribution (i.e., rotten tomatoes movie reviews), suggesting that similarities among calibrated coefficients may indicate potential transferability among tasks. To explore this further, we transfer the context vectors and calibrated linear coefficients from a source task to various target tasks. We then measure the differences between the transferred results and their respective zero-shot outcomes to identify both positive and negative transfers. Figure 4 (middle) shows that performance on MR can be effectively enhanced by transferring context vectors and calibrated coefficients from SST-2, and vice versa. Similarly, both SST-2 and MR benefit from transfers from SST-5, likely due to their shared focus on sentiment analysis. An intriguing aspect of our findings is the asymmetric nature of transferability among different tasks; a successful transfer in one direction does not necessarily guarantee success in the opposite direction (e.g., transfer between HateSpeech18 and SST-2).
@@
Enhance New Task with Existing Anchors I2CL posits two pivotal features distinct from typical ICL: (1) An interface that facilitates vector arithmetic; (2) A set of linear coefficients that act as “task-ids”, laying a foundation for effective transfer learning [Zhuang et al., 2021]. Let ${\bm{c}}_{1},\ldots,{\bm{c}}_{N}$ represent calibrated coefficients for existing tasks, and ${\bm{c}}_{\text{new}}$ for the new task. Associated with these coefficients, respectively, are context vectors ${\bm{v}}_{1},\ldots,{\bm{v}}_{N}$, and ${\bm{v}}_{\text{new}}$.
@@
where $\tau$ is the temperature. Finally, we reinitialize ${\bm{v}}_{\text{new}}$ and ${\bm{c}}_{\text{new}}$ using the weighted average of retained context vectors and coefficients: ${\bm{v}}_{\text{avg}}=\sum_{i\in I}P(i){\bm{v}}_{i},{\bm{c}}_{\text{avg}}=\sum_{i\in I}P(i){\bm{c}}_{i}$, and perform another around of calibration. We refer to Appendix B for detailed algorithms and implementation. Empirical results in Table 2 demonstrate a clear benefit of the proposed transfer learning method.
Injection Dynamics Thus far, we have demonstrated that a few static scalars, conditioned on appropriate context vectors, can successfully execute a wide range of NLP tasks. Herein, we delve into how context vectors are injected. Using the SST-2 as an example, we examine the coefficient values across different layers (additional plots in Appendix D). One reasonable speculation postulates a gradual addition of context vectors to the residual streams. Nevertheless, observations from Figure 4 (right) reveal a more nuanced control over the information injection process. Instead of monotonically adding ($+$) more information to the residual streams, I2CL also allows the deletion (--) of information at certain layers, with coefficient values fluctuating across different layers.
@@
Table 7: Vectorization method.
Name Accuracy (%) Zero-shot $58.13$ Concatenate $70.80$ PCA $75.22$ Average (ours) $75.12$

### v2

@@ 3.1 Benchmarking I2CL
Table 1: Comparison between I2CL and baseline methods on Llama2-7b. The best results are highlighted in bold, and the second-best results are underlined. In addition to a practical gauge of the inference speed and memory usage (see Fig 1), we include an examination of cached parameters. Here, $M$, $D$, and $L$ denote the number of demonstration tokens, model dimension, and architecture layers, respectively. $P$ indicates the number of extra learnable tokens in the Soft-prompt method, and $1/K$ represents the compression rate of corresponding context-compression method.
@@
Method SST-2 SST-5 TREC AGNews Subj HateSpeech18 DBPedia EmoC MR Avg. acc. ##\## cached (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ param. $\downarrow$ Zero-shot $83.00$ $27.00$ $50.00$ $70.20$ $51.40$ $54.20$ $72.00$ $41.80$ $73.60$ $58.13$ 00 Few-shot (ICL) $94.44$$\pm$1.44 $41.72$$\pm$3.68 $77.32$$\pm$4.41 $85.68$$\pm$2.00 $52.56$$\pm$3.09 $70.24$$\pm$5.80 $96.64$$\pm$0.48 $75.48$$\pm$1.63 $93.24$$\pm$0.50 $76.37$ $2MDL$ Noise vector $49.88$$\pm$0.24 $20.56$$\pm$0.64 $20.12$$\pm$10.92 $27.32$$\pm$2.82 $49.64$$\pm$0.48 $59.84$$\pm$8.04 $7.28$$\pm$0.37 $26.76$$\pm$3.04 $50.12$$\pm$0.24 $34.61$ $2DL$ Label-anchor $83.32$$\pm$5.95 $27.68$$\pm$4.21 $77.48$$\pm$3.49 $83.72$$\pm$1.04 $53.00$$\pm$2.95 $64.52$$\pm$8.09 $81.40$$\pm$3.67 $59.12$$\pm$10.60 $84.40$$\pm$5.89 $68.29$ $2(M/K)DL$ Task-vector $81.44$$\pm$4.73 $25.96$$\pm$0.59 $65.68$$\pm$1.93 $79.68$$\pm$4.07 $58.56$$\pm$4.91 $67.68$$\pm$3.70 $89.48$$\pm$2.58 $44.64$$\pm$3.53 $82.32$$\pm$5.37 $66.16$ $D$ ICV $86.28$$\pm$0.55 $33.48$$\pm$0.65 $63.84$$\pm$0.15 $72.40$$\pm$0.37 $56.56$$\pm$0.70 $60.56$$\pm$1.50 $73.64$$\pm$0.88 $49.16$$\pm$1.24 $84.04$$\pm$1.10 $64.44$ $DL$ I2CL (ours) $\mathbf{87.68}$$\pm$2.47 $\mathbf{39.12}$$\pm$2.69 $\mathbf{78.56}$$\pm$5.32 $\mathbf{85.48}$$\pm$1.16 $\mathbf{73.84}$$\pm$3.84 $\mathbf{69.88}$$\pm$5.67 $\mathbf{90.16}$$\pm$1.86 $\mathbf{63.72}$$\pm$1.37 $\mathbf{87.68}$$\pm$2.26 $\mathbf{75.12}$ $2DL$ AutoComp. $92.44$$\pm$3.29 $25.8$$\pm$4.8 $62.52$$\pm$9.34 $86.36$$\pm$1.03 $60.16$$\pm$0.32 $53.2$$\pm$6.1 $92.68$$\pm$2.86 $29.56$$\pm$5.07 $82.76$$\pm$7.34 $63.94$ $2(M/K)DL$ ICAE $91.64$$\pm$1.69 $38.8$$\pm$1.56 $50.92$$\pm$8.38 $80.48$$\pm$2.35 $50.52$$\pm$9.17 $65.48$$\pm$7.18 $62.08$$\pm$1.86 $54.04$$\pm$4.69 $89.48$$\pm$1.45 $64.83$ $2(M/K)DL$ CEPE $74.28$$\pm$3.9 $36.2$$\pm$0.56 $55.48$$\pm$3.42 $78.00$$\pm$3.49 $59.12$$\pm$1.6 $61.72$$\pm$5.26 $87.24$$\pm$1.2 $42.28$$\pm$3.31 $82.36$$\pm$1.61 $64.08$ $2(M/K)DL$
I2CL Achieves Few-shot Performance at Zero-shot Inference Cost  I2CL is designed to enhance zero-shot performance, providing an alternative for ICL under constrained scenarios. As shown in Table 1, I2CL significantly outperforms the zero-shot counterpart by 17% in absolute accuracy and is only marginally behind (by around 1% on average) the few-shot learning. Note that I2CL consumes only zero-shot inference cost in terms of both memory usage and inference speed. An interesting observation arises from the Subj task, where the instructions (i.e., demonstration examples) do not function effectively with ICL, yet are well adhered to under I2CL. We hypothesize that this phenomenon is due to the inherent properties of the pre-trained LLM, causing it to excel in certain tasks while lagging in others [^5] As demonstrated in Table 11, zero-shot performance of Subj is much higher under GPT-J than Llama2-7b though the later one is generally considered more powerful., and our proposed noisy self-calibration strategy can effectively rectify this deficiency.
Comparison with Inference-time Methods w/o Additional Data To further validate the efficacy of I2CL, we compare it with several comparable methods: (1) Noise vector: replacing the context vector with random noise to assess its necessity; (2) Label-anchor: a token reduction method from Wang et al. (2023), using formatting and label tokens as anchors; (3) Task-vector: task-vector (Hendel et al., 2023) also improves zero-shot performance without introducing additional computational and memory overhead at inference. It requires a hold-out validation set to identify the optimal replacement layer for each task at test-time; (4) ICV: we adapt In-context Vector Liu et al. (2024) method for our scenarios treating query tokens as negative examples and answer tokens as positive examples. We manually optimize ICV’s strength parameter for each task and report the best performance. Please refer to Appendix B for implementation details.
As demonstrated in Table 1, replacing the context vector with random noise significantly degrades performance, highlighting the critical role of context vector. Label-anchor method exhibits a decent upgrade over the zero-shot baseline, achieving results most comparable to I2CL. Nevertheless, its inference cost remains dependent on the length of demonstration tokens, with reductions proportional to their compression rate. Like I2CL, the task-vector method also enjoys zero-shot inference expenses; yet its performance is sensitive to the downstream task and is, on average, slightly inferior to the label-anchor method. Finally, a clear improvement over zero-shot baseline can be seen by adapting ICV method for few-shot classification tasks. Notwithstanding the boost, it necessitates manual effort for hyperparameter selection, which limits its applicability for scenarios that favor automated pipelines. Compared to the above methods, I2CL achieves the best performance on all tasks with neither manual intervention nor task-specific hyperparameter selection.
Figure 3: Scaling trend of I2CL.
Comparison with Inference-time Prompt Compression Methods Another line of work that leverages external large-scale datasets to learn an amortized context compressor can also be applied to ICL scenarios. To this end, we compare I2CL with three SoTA prompt compression methods: AutoCompressors Chevalier et al. (2023), ICAE Ge et al. (2023), and CEPE Yen et al. (2024). Here, we directly apply their released pre-trained models on our tasks to avoid implementation bias. As shown in the bottom section of Table 1, all three methods yield a decent improvements over zero-shot baseline, demonstrating their effectiveness in compressing context tokens. However, these methods produce mixed and fluctuating results across different tasks, likely due to their different training data and strategies. Instead of learning a universal compressor, I2CL opts for customizing a task-specific compression strategy at test time, rendering it additional adaptability for few-shot learning tasks. Nevertheless, we do not claim overall superiority of I2CL over prompt-compression methods as they are deliberately tailored for excessively long context compression which is orthogonal to the primary application of I2CL.
Table 2: Comparison between different PEFT-based few-shot fine-tuning strategies.
Method ##\## trainable SST-2 SST-5 TREC AGNews Subj HateSpeech18 DBPedia EmoC MR Avg. acc. params. (K) $\downarrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ (%) $\uparrow$ Prompt-tuning $4.10$ $56.24$$\pm$6.99 $24.24$$\pm$2.96 $55.20$$\pm$4.14 $78.00$$\pm$7.60 $57.40$$\pm$4.93 $49.56$$\pm$6.96 $74.40$$\pm$6.43 $35.08$$\pm$5.29 $54.32$$\pm$1.76 $54.94$ LoRA $4194.30$ $84.80$$\pm$6.59 $39.87$$\pm$4.33 $75.97$$\pm$10.77 $83.80$$\pm$2.32 $70.47$$\pm$10.68 75.32$\pm$2.88 $91.40$$\pm$3.54 $53.67$$\pm$16.27 $83.07$$\pm$0.25 $73.15$ IA3 $262.14$ 89.40$\pm$2.08 46.93$\pm$0.81 $75.41$$\pm$4.94 $84.43$$\pm$1.45 $56.67$$\pm$3.07 $62.54$$\pm$5.58 93.91$\pm$0.49 $59.75$$\pm$3.67 88.00$\pm$1.88 $73.00$ I2CL (ours) $\mathbf{0.13}$ $87.68$$\pm$2.47 $39.12$$\pm$2.69 78.56$\pm$5.32 85.48$\pm$1.16 73.84$\pm$3.84 $69.88$$\pm$5.67 $90.16$$\pm$1.86 63.72$\pm$1.37 $87.68$$\pm$2.26 $\mathbf{75.12}$
Comparison with Test-time PEFT Methods. The design of I2CL involves a lightweight training process (i.e., self-calibration). As such, we further establish a comparison between I2CL and three representative PEFT methods: prompt-tuning (Lester et al., 2021), LoRA (Hu et al., 2021), and IA3 (Liu et al., 2022) to underscore the effectiveness and efficiency of I2CL. As shown in Table 2, prompt-tuning fails under such data scarcity, while LoRA and IA3 perform on par with I2CL. Importantly, I2CL achieves the best overall performance with approximately 100x fewer learnable parameters than prompt-tuning, 1,000x fewer than IA3, and 10,000x fewer than LoRA. We refer readers to Appendix B for implementation details.
Scaling Property of I2CL Although I2CL is primarily designed for few-shot scenarios, it exhibits strong scaling properties. As shown in Fig. 3, ICL peaks at around 5-shots and additional demonstration examples do not necessarily bring further benefits. In contrast, I2CL can easily scale up to hundreds or even thousands of demonstration examples without performance degradation, and it readily surpasses few-shot performance under a modest number of demonstration examples. We note that using more demonstration examples increases the calibration cost.
@@ 3.2 On the Formation of the Context Vector
Deficient ICL $\bm{\neq}$ Deficient I2CL It is well-known that ICL is sensitive to the choice of demonstration examples (Dong et al., 2023), even their order (Zhao et al., 2021; Lu et al., 2022). In this context, we investigate whether poorly performing demonstration examples in ICL will similarly affect I2CL. We sample $20$ groups of demonstrations with random instances and orders, and identify the group with the poorest ICL performance on a hold-out dataset that is non-overlapping with both train and evaluation sets. Using these same demonstration examples, we perform both ICL and I2CL on the evaluation set. Referencing Figure 4 (left), the poorly performing demonstrations lead to a severely degraded ICL performance ($-7\%$), while I2CL performance remains largely unaffected ($-0.5\%$). This result suggests that I2CL requires less careful curation of demonstration examples than ICL. We hypothesize that this robustness stems from I2CL’s ability to extract task-relevant features that are invariant to surface-level variations in the demonstrations. To further corroborate this hypothesis, we visualize the context vectors using t-SNE (van der Maaten & Hinton, 2008) in Figure 4 (right). The visualization reveals that context vectors from different demonstration groups cluster tightly by task, indicating that I2CL captures consistent task-specific representations despite variations in the demonstrations.
Figure 4: Left: Evaluation of I2CL and few-shot learning under deficient demonstrations. The symbol * denotes the results under deficient demonstration examples. “Unseen demo” refers to the evaluation of calibrated coefficients on unseen demonstrations. Middle: Analysis of the influencing factors of context vectors. “Random-label” indicates random input-label mappings. “Random-order” refers to the random permutation of words. “W/o format” signifies excluding the template tokens during the creation of context vectors. Right: t-SNE plot of context vectors. Each circle denotes a context vector generated using a group of randomly sampled demonstration examples.
@@
Influencing Factors of the Context Vector Generation We have shown above that context vectors are robust against the variations in demonstration examples. Here, we explore various factors that may impact the context vector’s functionality. Inspired by counter-intuitive findings reported in (Zhang et al., 2022; Min et al., 2022b), we experiment with two variations of demonstration examples. Random-label: pairing each input sentence with a random label from the task, rather than the ground-truth label. Random-token: randomly permuting all words within a demonstration example. Another ineligible factor involves the inclusion of formatting tokens. To this end, we also evaluate on W/o-format: removing formatting words from a demonstration example, retaining only the input sentence and its corresponding label, e.g., “Review: This is a great movie. Label: positive. As revealed in Figure 4 (middle), input-label mapping relations have minimal impact on context vector formation, mirroring the observations made in ICL. However, these mapping relations are crucial for calibration purposes; using random input-label mapping during calibration undermines the functionality of I2CL. Unlike the phenomenon observed in (Zhang et al., 2022), word sequence holds essential statistics under a causal architecture, and randomly permuting input words yields degenerated context vectors, leading to a clear performance drop. Lastly, formatting tokens contribute substantially—removing them results in a noticeable performance degradation.
@@ 3.3 Analysis of Calibrated Linear Coefficients
With a grasp on the formation of context vectors, we delve into the properties of calibrated linear coefficients to enhance our understanding of the internal mechanisms of I2CL.
@@
Linear Coefficients Are Generalizable Given the demonstration-dependent context vectorization and calibration process, it is natural to consider I2CL as an inference-time optimization framework. However, we demonstrate that the coefficients are generalizable and require calibration only once per task. Concretely, we evaluate a set of calibrated coefficients on five new context vectors generated using five unseen groups of demonstrations. As exhibited in Figure 4 (left), the calibrated linear coefficients generalize well to unseen demonstrations. We attribute this generalization capability to two factors: (1) the inherent robustness of the context vector against token space variations, as validated in Section 3.2, and (2) the sufficiency of calibrated linear coefficients, independent of context vectors and query activation, to uniquely represent a task. We substantiate the second premise in the following subsection.
@@
Calibrated Coefficients Embed Task Semantics Here, we investigate whether calibrated coefficients alone are adequate to serve as task-ids. In pursuit of this goal, we concatenate calibrated coefficients across layers to form a one-dimensional vector, which we then visualize using t-SNE. As illustrated in Figure 5 (left), the calibrated linear coefficients are tightly clustered for instances associated with the same task and fall apart otherwise, providing complementary evidence for their good generalization capability. One exception comes from the proximity between MR and SST-2. This phenomenon is not unexpected, as both tasks originate from the same underlying distribution (i.e., rotten tomatoes movie reviews), suggesting that similarities among calibrated coefficients may indicate potential transferability among tasks. To explore this further, we transfer the context vectors and calibrated linear coefficients from a source task to various target tasks. We then measure the differences between the transferred results and their respective zero-shot outcomes to identify both positive and negative transfers. Figure 5 (middle) shows that performance on MR can be effectively enhanced by transferring context vectors and calibrated coefficients from SST-2, and vice versa. Similarly, both SST-2 and MR benefit from transfers from SST-5, likely due to their shared focus on sentiment analysis. An intriguing aspect of our findings is the asymmetry of transferability among different tasks; a successful transfer in one direction does not necessarily forecast the opposite (e.g., transfer between HateSpeech18 and SST-2).
@@
Enhance New Task with Existing Anchors I2CL posits two pivotal features distinct from typical ICL: (1) An interface that facilitates vector arithmetic; (2) A set of linear coefficients that act as task-ids, laying a foundation for transfer learning (Zhuang et al., 2021). Here, we present an transfer learning algorithm based on these features. Let ${\bm{c}}_{1},\ldots,{\bm{c}}_{N}$ represent calibrated coefficients for existing tasks, and ${\bm{c}}_{\text{new}}$ for the new task. Associated with these coefficients, respectively, are context vectors ${\bm{v}}_{1},\ldots,{\bm{v}}_{N}$, and ${\bm{v}}_{\text{new}}$.
@@
where $\tau$ is the temperature. Finally, we reinitialize ${\bm{v}}_{\text{new}}$ and ${\bm{c}}_{\text{new}}$ using the weighted average of retained context vectors and coefficients: ${\bm{v}}_{\text{avg}}=\sum_{i\in I}P(i){\bm{v}}_{i},{\bm{c}}_{\text{avg}}=\sum_{i\in I}P(i){\bm{c}}_{i}$, and perform another round of calibration. We refer to Appendix B for detailed algorithms and implementation. Empirical results in Table 3 demonstrate a clear benefit of the proposed transfer learning method.
Injection Dynamics Thus far, we have demonstrated that a few static scalars, conditioned on appropriate context vectors, can successfully execute a wide range of tasks. Herein, we delve into how context vectors are injected. Using the AGNews as an example, we examine the coefficient values across different layers (additional plots in Appendix C.4). One reasonable speculation postulates a gradual addition of context vectors to the residual streams. Nevertheless, observations from Figure 5 (right) reveal a more nuanced control over the information injection process. Instead of monotonically adding ($+$) more information to the residual streams, I2CL also allows the deletion (--) of information at certain layers, with coefficient values fluctuating across different layers.